To Log in Use for check in bed * Il CorComm Cole New W owowe Chew O w + For 1 () the n the re
Answer #1
20y3 20y2
Accounts Receivable turnover ratio 8.2 7.5
No of days of sales in receivables: 44.51 48.67

The collection of accounts receivable has INCREASED. This can be seen in both INCREASE in Accounts Receivable Turnover and DECREASE in collection period.

Notes:

Accounts receivable turnover ratio is obtained by dividing net credit sales by average accounts receivable.

So,

  • 20y3 = 5637500/ [(725000+650000)/2] = 8.2 Times
  • 20y2 = 4687500/ [(650000+600000)/2] = 7.5 Times

No of days of sales in receivables:

  • 20y3 = 365/8.2 = 44.51 days
  • 20y2 = 365/7.5 = 48.67 days

(Please comment in case of any query regarding the solution)